OpenStudy (anonymous):

Dorothy Wagner is currently selling 50 "I ♥ Calculus" T-shirts per day, but sales are dropping at a rate of 9 per day. She is currently charging $7 per T-shirt, but to compensate for dwindling sales, she is increasing the unit price by $1 per day. How fast, and in what direction, is her daily revenue currently changing?

OpenStudy (anonymous):

I know that it's decreasing, and I got by $18 a day, but apparently I'm wrong

OpenStudy (anonymous):

50*7 = 350 41*7 = 287 350-287 =63 increasing the $1 50*8 = 400 42*8 = 336 400-336 = 64